# Safe MPC Alignment with Human Directional Feedback

### v2

@@ VIII-C Result and Analysis
In the real-world robot experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$. The results show the efficacy and efficiency of our method in real-world robot learning applications. To illustrate the robot learning progress, in Fig. 22, we use one trial to show the robot behavior at different numbers of human corrections.
@@
(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, including both the human aspects and hardware aspects. On human aspects, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations. On hardware side, noisy correction signals and the possible timing delay between the recorded human correction and robot state can contribute to the failure of the learning.
(II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons. First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game Appendix B-D. This makes the safety constraints easy to learn for the real-world task. Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in real world. A good perception of the robot’s state could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
VIII-D Discussion
Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
VIII-D1 Learning Task-Independent Safety Constraints
In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint could be task-specific, instead of task-independent. For instance, in the drone navigation game, since the three task targets are placed above the gate, the learned constraint shown in Fig. 19 can only reflect the upper edge of the unknown gate. One limitation of this is that the learned task constraints have poor generalization to new tasks. For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig. 19 will fail.
To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples. By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints reusable in new tasks.
VIII-D2 Robustness Against Human Mistakes
@@
Figure 23: One illustration of the failure case in the drone game user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
@@
Another is to add a selection mechanism on human correction. For example, one can use some statistical method to rule out mistake corrections before performing a hypothesis cut. We will leave those possible extensions in our future work.
IX Conclusion
This paper proposes safe MPC alignment method to online learn a safety constraint in safe MPC controllers from human directional feedback. The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback. The method has a provable guarantee for the total number of human feedback needed for learning, and also can declare hypothesis misspecification. Numeric examples, user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness of the proposed approach. The results show that the proposed method can learn safety constraints effectively and efficiently. To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
@@ A-A Proof of Lemma 2
We prove this lemma by induction. First, we have $\bar{\Theta}_{H}\subset\Theta_{0}$. Suppose $\bar{\Theta}_{H}\subset\Theta_{i-1}$, $i=1,2,3,...$. Due to the condition $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$ in (12), following the update (13), we have $\bar{\Theta}_{H}\subset\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$, which completes the proof. ∎
@@ A-B Proof of Lemma 3
(14) can be directly obtained from (12), (2) and (5). Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
$\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (37)
with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
Multiplying the denominator on both side and substituting the weight-feature expression of $g_{\boldsymbol{\theta}}$ in (5) yields $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}$ in Lemma 3. The condition $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields $\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
@@ A-C Proof of Lemma 4
$\mathbf{a},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})\right>=0,$ (39)
@@
$\mathbf{a},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>-\gamma\frac{1}{g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{a},\nabla g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>=0.$ (40)
Due to $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$, one can multiple $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ on both sides of the above equality, which leads to $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$ based on the definition (15). The second term in (18) directly follows $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. This concludes the proof. ∎
@@ A-D Proof of Lemma 5
By Lemma 4, the MVE center $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ is on the cutting hyperplane $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$. Directly from [63], one has
@@ A-E Proof of Theorem 1
We complete the proof in two parts. First, we prove the number of learning iterations is less than $K$, then we use this to prove there will be a $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$, $k=0,1,2...,K-1.$
@@ A-F Proof of Theorem 2
We first prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$. Given the condition $\Theta_{0}\cap\bar{\Theta}_{H}=\emptyset$, the algorithm will never terminate because $\forall i=1,2,3,...\infty$, $\mathcal{C}_{i}\supset\bar{\Theta}_{H}$ due to Lemma 2. Since the volume of the initial hypothesis space is finite, by applying Lemma 5 we have
@@ B-A Inverted Pendulum
When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $[0,0]^{\mkern-1.5mu\mathsf{T}}$ and $[\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$. Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
B-B Two-link Robot Arm
We assume each links has unit length and mass, and the mass is evenly distributed along the link. The corresponding inertia matrix and Coriolis term can then be calculated.
B-C Drone Navigation Game
@@ B-C1 Environment
The drone starts from $[0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. We set three different targets at positions $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. For the narrow gate (yellow rectangle) in Fig. 15, both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate. In the drone MPC policy (1), the dynamics and its parameters are set following ([20]); and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
@@ B-C2 Human Correction Interface
The motion of the drone is rendered in two camera views in MuJoCo visualization (Fig. 15) for better human perception: one is the first-person view tracking the drone, the other a third-person fixed view. The gate, the current MPC planning solution, and the target position are visible to the users.
Human users apply corrections to the drone via a keyboard during the drone flying. The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table V. The safety constraint (54) in the drone’s MPC policy is immediately updated once a user correction is detected. When the user presses "ENTER", the drone is reset to the initial position and target will randmize.
@@
"up" $[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
"down" $[1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
"left" $[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
"right" $[0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
"shift" $[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
"ctrl" $-[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)

### v3

@@ VIII-C Result and Analysis
In the real-world experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$. The results show the efficacy and efficiency of our method in real-world applications. To illustrate the learning progress, Fig. 22 shows the robot behavior after different numbers of corrections in one trial.
@@
(I) Real-world experiment has slightly lower success rate: While the difference between simulation and real-world studies is small in success rate, we have identified some factors for such a difference. On human aspects, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations. On hardware side, noisy correction signals and the possible timing delay between the recorded correction and the robot state can contribute to failure.
(II) Real-world experiment needs less human corrections: This could be attributed to two reasons. First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game. This makes the safety constraints easy to learn for the real-world task. Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in the real world. A good perception of the robot’s state could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
IX Discussion and Conclusion
IX-A Discussion
Based on the above hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
IX-A1 Learning Task-Independent Safety Constraints
In the real-world and simulation experiments, the learned safety constraint could be task-specific, instead of task-independent. For instance, in the drone navigation game, since the three task targets are placed above the gate, the learned constraint shown in Fig. 19 can only reflect the upper edge of the unknown gate. One limitation is that the learned constraints have poor generalization to new tasks. For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig. 19 will fail.
To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples. By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region.
IX-A2 Robustness Against Human Mistakes
@@
Figure 23: A failure case in the drone game user study. A user makes a wrong "up" correction at the beginning and fails to correct the motion of the drone.
@@
Another is to add a selection mechanism for human correction. For example, one can use a voting system to preserve more fraction of the hypothesis space and avoid cutting out the human safety set, as shown in [xie2025robust].
IX-A3 Limitation of Weight-Feature Constraints
In this work, we assume the set of task-relevant features in $\boldsymbol{\phi}(\boldsymbol{\xi})$ is selected as a priori. These features define the space in which the safety constraints are represented. However, in practice, identifying appropriate features can be challenging. Poorly chosen features may fail to capture the intended safety constraints, leading to undesired learning outcomes. Furthermore, feature representations are often task-dependent—features that are effective in one task may not generalize to another. This task-specific dependence inherently limits the development of a unified framework for constraint learning across diverse tasks, making the lack of task-generalizable feature selection a key limitation of the proposed method. At the same time, learning unstructured constraints entirely from scratch would demand extensive human interaction and involvement, rendering such approaches impractical for most real-world applications. This challenge likely explains why recent works on learning from human feedback [bajcsy2017learning, jin2022learning, losey2022physical] commonly assume task-specific feature structures or prior knowledge of the environment/tasks.
We recognize that identifying or learning a universal feature representation that can flexibly describe a wide variety of safety constraints remains an important and open research challenge. We consider this an exciting direction for future research, as it could significantly increase the applicability and effectiveness of safety constraint learning in real world.
IX-B Conclusion
This paper proposes a Safe MPC Alignment method to online learn a safety constraint in Safe MPC controllers from human directional feedback. The method, named hypothesis space cutting, has a geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback. To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback. The method has a provable guarantee for the total number of human feedback needed for learning, and also can declare hypothesis misspecification. Numeric examples, user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness of the proposed approach. The results show that the proposed method can learn safety constraints effectively and efficiently.
@@ A-A Proof of Lemma 2
We prove this lemma by induction. First, we have $\bar{\Theta}_{H}\subset\Theta_{0}$. Suppose $\bar{\Theta}_{H}\subset\Theta_{i-1}$, $i=1,2,3,$.... From the previous assumptions (7) and (9), $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$. Following the update (13), we have $\bar{\Theta}_{H}\subset\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$, which completes the proof. ∎
@@ A-B Proof of Lemma 3
(14) can be directly obtained from (12), (2) and (5). Based on the definition of the penalty objective function (2), $\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
$\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (36)
with a positive denominator $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
Multiplying the denominator on both side and substituting the weight-feature expression of $g_{\boldsymbol{\theta}}$ in (5) yields $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}$ in Lemma 3. The condition $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields ${\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
@@ A-C Proof of Lemma 4
$\left<\mathbf{a},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})\right>=0,$ (38)
@@
$\left<\mathbf{a},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>-\gamma\frac{1}{g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{a},\nabla g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>=0.$ (39)
Due to $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$, one can multiply $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ on both sides of the above equality, which leads to $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$ based on the definition (15). The second term in (18) directly follows $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. This concludes the proof. ∎
@@ A-D Proof of Lemma 5
By Lemma 4, the MVE center $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ is on the cutting hyperplane $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$. Directly from [tarasov1988method], one has
@@ A-E Proof of Theorem 1
We complete the proof in two parts. First, we prove the number of learning iterations is less than $K$, then we use this to prove there will be a $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$, k=0,1,2,K1.k=0,1,2...,K-1.
@@ A-F Proof of Theorem 2
We first prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$. Given the condition $\Theta_{0}\cap\bar{\Theta}_{H}=\emptyset$, the algorithm will never terminate because $\forall i=1,2,3,$$\infty$, $\mathcal{C}_{i}\supset\bar{\Theta}_{H}$ due to Lemma 2. Since the volume of the initial hypothesis space is finite, by applying Lemma 5 we have
@@ B-A Inverted Pendulum
When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $0,0]^{\mkern-1.5mu\mathsf{T}}$ and $\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$. Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
B-B Simulated Quadrotor Navigation
B-B1 Environment
The quadrotor starts from $0.0,0.0,5.0]^{\mkern-1.5mu\mathsf{T}} [$m]. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. The target is to navigate to the target position $15.0,0.0,10.0]^{\mkern-1.5mu\mathsf{T}}[$m] within a zigzag tube. The average radius of the tube section is set to be 2.5m.
B-B2 Learning Neural Features
We construct the neural feature $\boldsymbol{\phi}$ using a neural Signed Distance Function (SDF) model trained on synthetic 3D point samples around the tube surface. Each input coordinate (scaled for stability) is first encoded using sinusoidal positional embeddings, then passed through a 5-layer MLP (width 128, ReLU activations). The positional embeddings are skip-connected to the input of the 5th hidden layer. The penultimate layer outputs a 16D feature vector used as $\boldsymbol{\phi}$; the final layer predicts a scalar signed distance supervised during training.
Training data is generated by randomly perturbing surface points on the tube. Signed distance labels are computed as the offset from the nearest point on the tube surface to the sample point, and normals are estimated from the same displacement vector pointing from the nearest surface point to the sample point. The model is trained with three loss terms: $L_{2}$ distance loss, cosine similarity loss on normals (via auto-diff), and Eikonal loss to regularize gradient norm, which is similar to the work [ortiz2022isdf]. Optimization is done using Adam for 250 epochs with a batch size of 1024.
B-B3 Simulated Correction and Stop Behavior
Correction signals $\boldsymbol{a}_{x}=[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}},\boldsymbol{a}_{y}=[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}},\boldsymbol{a}_{z}=[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ in drone control space is used throughout the experiments, corresponding to different motion directions in cartesian space. During the process of learning positional constraints with polynomial features or neural features, if the distance between the quadrotor and the nearest point on the surface of the tube is less than 0.6, a simulated correction $\boldsymbol{a}$ is applied in the action space of the drone, based on the vector $\boldsymbol{d}=(d_{x},d_{y},d_{z})$ pointing from the closest tube surface point to the drone position:
$\boldsymbol{a}=d_{y}\boldsymbol{a}_{y}+d_{z}\boldsymbol{a}_{z}.$ (53)
Intuitively, the simulated correction "pulls" the drone to the center of the tube (designated safe region). After each correction, the navigation restarts, emulating an emergency stop. During the process of learning velocity constraint, Simulated human corrections against the velocity direction of form:
$\boldsymbol{a}=-v_{x}\boldsymbol{a}_{x}-v_{y}\boldsymbol{a}_{y}-v_{z}\boldsymbol{a}_{z}$ (54)
are applied whenever the velocity $\boldsymbol{v}=(v_{x},v_{y},v_{z})$ exceeds 0.4, and an emergency stop is triggered if it surpasses 0.45.
B-C Quadrotor Navigation Game with Real Human Feedback
@@ B-C1 Environment
The drone starts from $0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. We set three different targets at positions $19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. For the narrow gate (yellow rectangle) in Fig. 16(a), both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate. In the drone MPC policy (1), the dynamics and its parameters are set following [jin2022learning]; and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
@@ B-C2 Human Correction Interface
The drone’s motion is visualized in MuJoCo using two complementary camera views: a first-person tracking view and a fixed third-person view. The gate, the current MPC plan, and the target position are always shown to the user. Human users provide corrections through keyboard inputs during flight. Key–command mappings are given in Table V. Whenever a correction is detected, the safety constraint (55) in the MPC policy is updated instantly. Pressing ENTER resets the drone to its initial state and reset the target.
@@
"up" $1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
"down" $1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
"left" $0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
"right" $0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
"shift" $1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
"ctrl" $1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)